Consider the discrete-time LTI system with impulse response
h(n) = n u(n).
(a) Find and sketch the output y(n) when the input x(n) is
x(n) = δ(n) - 2δ(n - 5) + δ(n - 10).
1
Expert's answer
2020-10-29T08:23:34-0400

"h(n) = nu(n)"

"x(n) = \delta(n)-2\delta(n-5)+\delta(n-10)"


Output "y(n)" is convolution of "h(n)" and "x(n)" .i.e


"y(n) = x(n) * h(n)" (here '"*"' implies convolution)


i.e. "y(n) = \displaystyle\sum_{k=-\infin}^\infin" "x(n)h(n-k)"


i.e. "y(n) = \displaystyle\sum_{k=-\infin}^\infin" "[\delta(k)-2\delta(k-5)+\delta(k-10)](n-k)u(n-k)"


"y(n) = \displaystyle\sum_{k=-\infin}^\infin [(n-k)u(n-k)\delta(k) - 2(n-k)u(n-k)\delta(k-5)+\\(n-k)u(n-k)\delta(k-10)]"

This equation on doing summation we get,


"y(n) = nu(n)-2(n-5)u(n-5)+(n-10)u(n-10)"


Substituting values for n we get ,


"y(n) ={0,1,2,3,4,5,6,4,3,2,1}"


i.e "y(0) = 0, y(1)=1,y(2) = 2. \space so\space on \space and\space y(10) = 1"


i.e for all other values of n other than "0\le n \le 10" we have "y(n)=0"


The output could be plotted as shown in the attached image.